The solubility product of solid silver(I) bromide is 7.7 x 10−13 in water at 298 K. What is its solubility in grams per liter?

A. 1.7 x 10−4 g/L

B. 8.8 x 10−7 g/L

C. 1.8 x 10−7 g/L

D.1.7 x 10−5 g/L

E. 2.2 x 10−20 g/L
Answer #1

Let s be the solubility of AgBr.

Then, Ksp = [Ag+][Br-]

7.7 *10^(-13) = s*s

s = 8.775*10^(-7) M

Molar mass of AgBr = 187.77 g/mol

Then, s = 8.775*10^(-7)*187.77

s = 1.65*10^(-4) = 1.7*10^(-4) g/L.

Therefore, option (A) is the correct answer.

Please give good rating.